OpenStudy (anonymous):

Solve this application using logarithms. At what interest rate (to the nearest hundredth of a percent) compounded annually will money in savings double in five years?

OpenStudy (elonasushchik):

(n) = I(o)(1 + r/100)^n is the formula for compound interest. 1). 2 = 1(1 + r/100)^5 2 = (1 + r/100)^5 2^(1/5) = 1 + r/100) 0.2log(10)2 = (100 + r) / 100 = log(10)(100 + r ) - log(10)100 0.060206 = log(10)(!00 + r) - 2 2.060206 = log(10)(100 + r) Antilog . 114.8698 = 100 + r 14.6898 = r r = 14.87%

OpenStudy (elonasushchik):

Medal and fan?! Thx